### v1

# Is it possible to describe an electron by the evolution of a single point?

Abstract
The answer to the title-question is positive. The analysis of the geometry of continuous and differentiable curves in three-dimensional Euclidean space suggests that the point represents the location of the center of charge of the electron, satisfies a system of ordinary differential equations of fourth order, and moves at the speed of light. The center of mass of the electron is a different point and will be determined by the evolution of the center of charge. It is the relative motion of the center of charge around the center of mass that gives rise to the spin and magnetic properties.
1 Introduction
Let us consider the possibility of describing a classical spinning electron by the evolution of a single point $\bi r$} in three-dimensional Euclidean space. From the physical point of view it represents a mechanical system of three degrees of freedom. If we achieve this, it means that all observables related to the structure of the electron, like energy, linear momentum, spin, magnetic moment, electric dipole moment, etc., can be expressed in terms of the point $\bi r$} and its subsequent time derivatives.
The theory of continuous and differentiable curves in three-dimensional space was developped in 1847 and is known as the Frenet-Serret formalism. The points of the curve are parameterized by the arc length $s$ that represents the Euclidean distance along the curve of this point ${\bi r}(s)$ from some initial position ${\bi r}(0)$. The function ${\bi r}(s)$ and its subsequent derivatives with respect to the parameter $s$ are continuous and differentiable functions of this parameter.
Along the curve, and in terms of the derivatives of the point ${\bi r}(s)$ up to the third order, it is defined a system of three orthogonal unit vectors attached to the point. They are called the tangent $\bi t$}, the normal $\bi n$} and the binormal $\bi b$}. The characteristic of any curve are the two continuous and differentiable scalar functions $\kappa(s)$ and $\tau(s)$, known as the curvature and torsion, respectively, defined at every point ${\bi r}(s)$. The curvature is expressed in terms of the first two derivatives of the function ${\bi r}(s)$ and the torsion involves the derivatives up to the third order. The knowledge of the two continuous functions $\kappa(s)$ and $\tau(s)$, for $s\in[0,L]$, together the boundary conditions ${\bi r}(0)$ and ${\bi t}(0)$, ${\bi n}(0)$ and ${\bi b}(0)$, completely determine the curve of length $L$, in three dimensional space.
In three-dimensional Euclidean space, given any arbitrary but continuous and differentiable curve ${\bi r}(s)$, the derivatives $d{\bi r}(s)/ds$, $d^{2}{\bi r}(s)/ds^{2}$ and $d^{3}{\bi r}(s)/ds^{3}$, will be in general, three linearly independent vectors. If we take the next order derivative $d^{4}{\bi r}(s)/ds^{4}$, it can be expressed as a linear combination of the other three derivatives where the coefficients are functions of the four derivatives. This means that in the Frenet-Serret formalism the most general differential equation satisfied by a continuous and differentiable geometric curve in three dimensional space is a system of differential equations of fourth order where the independent variable is the arc length $s$.
Let us go one step forward. Instead of doing Geometry let us make Physics by introducing clocks. Let us consider that we describe the electron by the time evolution in three-dimensional space of a point ${\bi r}(s)$. In the above geometrical analysis it will mean that there exists a continuous and differentiable function $s(t)$ between the arc length $s$ and the elapsed time $t$, that will represent from now on the parameter of the curve ${\bi r}(t)\equiv{\bi r}(s(t))$. Let us consider that an inertial observer describes the free motion of this electron. Because in general, all derivatives are different from zero, the motion of the point is accelerated and therefore this point ${\bi r}(t)$ cannot represent the location of the center of mass (CM) of the electron. It can represent perhaps the location of another property of the electron.
If we are able to obtain a Lagrangian description of this system of only three degrees of freedom, the free Lagrangian $L_{0}$ has to depend up to the second time derivative of the point $\bi r$}, to obtain from the Euler-Lagrange equations a system of differential equations of fourth order. If this electron interacts there will be an interaction Lagrangian $L_{I}$, function of the interacting parameter, the charge, and the point $\bi r$} and some of its time derivatives. This is telling us that the point $\bi r$} represents the point where the external force is defined, i.e., the center of charge (CC) of the electron. This point has to be a different point than the possible CM, because in the free motion the CM must describe a straight line with no acceleration.
If we are able to describe an electron by a single point $\bi r$}, this point represents the evolution of the CC of the electron. We are not assuming any size or shape for the electron. We consider only that we have a characteristic point $\bi r$} that represents the center of charge where the external forces are defined, and that satisfies a system of fourth order differential equations.
Let us consider that in an inertial reference frame the motion is free. This means that for this inertial observer the displacement of the point at any time $t$, $ds(t)$ must be independent of the time. Otherwise if at two different times $t_{1}$ and $t_{2}$, $ds(t_{1})\neq ds(t_{2})$ this means that something different is happening at time $t_{2}$ than at time $t_{1}$, that produces a different displacement, contradictory with the idea that the motion is free. For this inertial observer the free motion of the electron is such that the CC is moving at a velocity of constant absolute value. If this description has to be independent of the inertial observer we arrive to the conclusion that we can never find an inertial observer at rest with respect to this point. If at some instant $t$, an inertial observer is at rest with respect to the point ${\bi r}(t)$, the velocity of this point is 0, but because this point is accelerated, at time $t+dt$ the velocity will be different from zero, contradictory with the condition that the velocity is of constant absolute value in any inertial frame. The velocity of the CC of an electron described by a single point must necessarily be an unreachable velocity of constant absolute value for all inertial observers.
This selects that the relationship between inertial observers must contain the condition of the existence of a velocity limit for the motion of the CC of the electron, and that the relative velocity among inertial observers can never reach this limit velocity. This limit velocity is an intrinsic property and must be the same for all inertial observers.
For the free motion also $\kappa(t)$ and $\tau(t)$, or its ratio $\kappa(t)/\tau(t)$ must be independent of the time, so that the trajectory of the CC is a helix. The center of the helix is a good candidate for the location of the CM that in the free motion is moving along a straight line at a constant velocity that can never reach the velocity of the CC. But the location of this point and the angular velocity of the comoving Frenet-Serret frame are completely determined by the point $\bi r$} and its subsequent time derivatives. The comoving frame rotates and the CC moves around the CM and because of this rotation and the orbital relative motion, this object has angular momentum. The CC has a trembling motion (zitterbewegung) around the CM. The particle described in this way has an internal frequency associated to this relative motion.
This simple analysis discards the use of the Galilei group as the group that defines the class of equivalent observers. If we consider that the three-dimensional space is infinite, the kinematical group of the possible formalism for describing an electron by a single point $\bi r$}, is the Poincaré group and the velocity of the CC is necessarily the speed of light.
2 General description and conclusions
The previous introduction is basically the Preamble of the Lecture Notes [1]. However, from a historical point of view, we arrived at this description of the electron by a different path. In 1985 [2] we analyzed the Lagrangian description of a mechanical system such that the boundary variables manifold that define the initial and final states of its variational evolution were considered to be a homogeneous space of the Galilei group. With this statement many models of non-relativistic spinning particles were described. The relativistic version of the above models were obtained in 1989 [3] by assuming that the boundary variables manifold of the variational formalism were homogeneous spaces of the Poincaré group. In that article we also assumed that the velocity of the point $\bi r$} could be the speed of light. Different models of relativistic spinning particles were found and also a spinning model of a classical photon described as a point and a comoving cartesian frame that rotates along the direction of the velocity. The spin of the photon has the same direction than the angular velocity, but it is an invariant mechanical property. The angular velocity and thus, the frequency of the photon, transform according to the Doppler effect transformation while the absolute value of the spin remains invariant.
This idea can be generalized to any space-time symmetry group of transformations. The kinematical group not only defines the space-time symmetries of the mechanical system. The homogeneous spaces of any Lie group are described by the same kind of parameters as the group elements, and, therefore, the set of the group parameters also supply and restrict the maximum number of classical variables to describe a classical elementary spinning particle.
The idea underlying this definition of a classical elementary particle is based on Wigner’s definition of an elementary particle in quantum formalism. According to Wigner the Hilbert space of the states of an elementary particle is a representation space of a unitary irreducible representation of the Poincaré group [4]. Irreducible means that the complete Hilbert space can be generated from any state. By selecting any arbitrary state, we can obtain all Poincaré transformations of this arbitrary state. Now let us produce all finite linear combinations of these new states and finally the closure of these states by including all limits of Cauchy sequences of this kind of vectors. We have generated the whole Hilbert space. This is the meaning that the representation is irreducible. This is a group theoretical definition of elementary particle.
The challenge is to find a definition based on group theory within the classical Lagrangian formalism. If we select any initial state of the variational description and transform this state by any Poincaré transformation we can obtain all possible physical states equivalent to the first one. If the mechanical system has no more classical states than the ones obtained by the above method, then, the set of states, i.e., the boundary variables manifold, is a homogenous space of the Poincaré group. If the mechanical system is not an elementary particle and we consider an excited state, we can never reach that state from the ground state by a simple change of the reference frame. The set of states of a non-elementary mechanical system is not a homogenenous space of the Poincaré group.
Later in 1994, we showed that all Lagrangian systems obtained in this way can be quantized, producing all known relativistic and non-relativistic quantum mechanical equations. In particular, the relativistic model that the CC is moving at the speed of light is the only model that satisfies Dirac’s equation when quantized [5]. All these results were collected in 2001 in the book [6].
The idea that leads to the definition of a classical elementary particle was revisited in 2008 in the form of a new fundamental principle, the Atomic principle [7]. The definition of elementary particle represents the physical fact that the elementary particle has no excited states and if it is not anihilated by some interaction its internal structure cannot be modified. A bound system like an atom can have excited states but an elementary particle cannot. Any mechanical system that contains some internal parts can be deformed but an elementary particle cannot. An elementary particle cannot be deformed and therefore all its possible states are just kinematical modifications of any one of them. In this way the whole classical theory of elementary spinning particles is based on three fundamental principles:
1. The Restricted Relativity Principle, requires that the inertial observers are related by the Poincaré group and the dynamical equations must remain invariant under the transformations of this group.
2. The Variational Principle, postulates the existence of a Lagrangian that describes the elementary particle such that the dynamical equations are the Euler-Lagrange equations, and with the above principle they must be Poincaré invariant.
3. The atomic Principle, i.e., the above definition of elementary particle. If the initial state in the Lagrangian description in an inertial reference frame is $x\equiv(x_{1},\ldots,x_{n})$, described by an unknown set of $n$ classical variables and this state changes at a later time to another $y\equiv(y_{1},\ldots,y_{n})$, and if the internal structure of the elementary particle has not been modified, then it is possible to find at this later time, another inertial observer, related to the previous one by means a Poincaré transformation $g$, such that the variables $y$ for this observer take exactly the same values as the variables $x$ for the previous observer. The the state $x$ is transformed into the $y$, such that the state y=gxy=g*x or $x=g^{-1}*y$. And this has to be valid for any pair of states. We recover that the boundary variables manifold of the Lagrangian description is a homogenous space of the Poincaré group.
The quantum mechanical theory of elementary particles is also based on three fundamental principles. We have to substitute the Variational Principle by Feynman’s Quantization Principle such that the Hilbert space that describes the quantum mechanical states is such that any wave function is a complex, squared integrable function defined on the boundary variables manifold of the corresponding classical Lagrangian description [5]-[6]. With this quantization method, under an arbitrary Poincaré transformation, the wave function $\psi(x)$ transforms with a unitary irreducible representation of the Poincaré group if the point $x$ belongs to a homogeneous space of the group. For other groups, like the Galilei group, what is obtained is that the wave function transforms with a projective unitary irreducible representation of the group, i.e., a unitary representation up to an arbitrary phase. This is a technical feature that depends on the structure of the group. If the group has nontrivial central extensions, like the Galilei group, then the representation is not a true representation but a projective representation. The Poincaré group has only trivial central extensions and thus all representations are true representations. In the Lagrangian formalism, Lagrangians are Poincaré invariant but not Galilei invariant. They transform with a time derivative of a gauge function [14].
The formalism can be extended to any kinematical group, Galilei, Poincaré, Weyl, Conformal or De Sitter groups, or any group we consider that describes the space-time symmetries of our mechanical system. If the rotation group is a subgroup of this general kinematical group we always obtain that the elementary particles described in this formalism have spin. If for instance, we select as the boundary variables manifold the space-time, i.e. the states are defined by the variables time $t$ and the position of a point $\bi r$}, $x\equiv(t,{\bi r})$, this manifold is in fact a homogeneous space of the Galilei or Poincaré group, and therefore it can be used to describe an elementary particle: The spinless point particle. But in Nature there are no spinless elementary particles. We need to include some more variables that transform under rotations, but we are restricted to a set that belongs to a homogeneous space of the kinematical group.
For the Galilei and Poincaré groups the maximum number of variables is 10, the number of parameters of the group. The largest homogeneous space is spanned by the variables $(t,{\bi r},{\bi u},\boldsymbol{\alpha})$, interpreted as the time $t$, the location of a point $\bi r$}, the velocity of this point ${\bi u}=d{\bi r}/dt$ and the orientation $\boldsymbol{\alpha}$, of a comoving Cartesian frame attached to the point $\bi r$}. The three parameters $\boldsymbol{\alpha}$ represent the 3 essential parameters that define a set of three orthogonal unit vectors ${\bi e}_{i}$, $i=1,2,3$, with origin at the point $\bi r$}, that describe the attached Cartesian frame. The constraint $u=c$ selects, in the Poincaré case, the only model that satisfies Dirac’s equation when quantized [5]. Because the Lagrangian has to depend on the next order time derivative of the above boundary variables, it is also a function of the acceleration of the point and the angular velocity of the attached comoving frame, and thus the point $\bi r$} satisfies a system of ordinary fourth-order differential equations. The spin observable is only related to the rotation invariance of the theory and is obtained in the Lagrangian formalism by applying Noether’s theorem to the invariance under rotations.
2.1 The Classical Dirac Particle
In addition to the previous references, we have published an article [8] where we analyze in detail what we call the classical Dirac particle, i.e., the relativistic model obtained by the general formalism that satisfies Dirac’s equation when quantized. All observables can be expressed in terms of the point $\bi r$} and its subsequent time derivatives. In particular, the center of mass (CM) $\bi q$} is defined with simultaneity with the CC, at every inertial frame, and it moves with a velocity ${\bi v}=d{\bi q}/dt$, such that $v<c$. The CC velocity ${\bi u}=d{\bi r}/dt$ always satisfies $u=c$. In terms of the CM velocity it is found that the energy and linear momentum can be written as
$H=\gamma(v)mc^{2},\quad{\bi p}=\gamma(v)m{\bi v},$ (1)
like in the case of the point particle, where the factor $\gamma(v)=(1-v^{2}/c^{2})^{-1/2}$. With the definition of the CM $\bi q$}
${\bi q}={\bi r}+\frac{c^{2}-{\bi v}\cdot{\bi u}}{a^{2}}{\bi a},\quad{\bi a}=\frac{d{\bi u}}{dt},$ (2)
where $\bi a$} is the acceleration of the CC, the fourth order differential equations of the point $\bi r$} can be separated into a coupled system of second order differential equations for both points $\bi q$} and $\bi r$}.
The dynamical equations of the center of mass $\bi q$} and the center of charge $\bi r$} of the Dirac particle under any electromagnetic field obtained from the general Lagrangian $L=L_{0}+L_{em}$, where $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ is the free Lagrangian and $L_{em}=-eA_{0}(t,{\bi r})+e{\bi u}\cdot{\bi A}(t,{\bi r})$, in terms of the scalar and vector potentials $A_{0}$ and $\bi A$} defined at time $t$ at the center of charge position $\bi r$}, respectively, are:
$\frac{d{\bi q}}{dt}$ $=$ ${\bi v},\quad\frac{d{\bi v}}{dt}=\frac{1}{m\gamma(v)}\left[{\bi F}-\frac{\bi v}{c^{2}}\left({\bi F}\cdot{\bi v}\right)\right],$ (3)
$\frac{d{\bi r}}{dt}$ $=$ ${\bi u},\quad\frac{d{\bi u}}{dt}=\frac{c^{2}-{\bi v}\cdot{\bi u}}{({\bi q}-{\bi r})^{2}}({\bi q}-{\bi r}),$ (4)
where the external force ${\bi F}=e{\bi E}(t,{\bi r})+e{\bi u}\times{\bi B}(t,{\bi r})$, is the Lorentz force, and $m$ and $e$ are the mass and the electric charge of the electron, respectively. The two velocities satisfy the constraints $u=c$, $v<c$. In the derivation of these dynamical equations we have made no use of a possible radiation reaction force. Equation (3) is just $d{\bi p}/dt={\bi F}$ and equation (4) is equation (2) when writing on the left hand side the CC acceleration $\bi a$}.
The free Lagrangian $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$, is translation invariant and it is an invariant function in terms of the velocity $\bi u$} and acceleration $\bi a$} of the CC, and also of the angular velocity $\boldsymbol{\omega}$ of the attached rotating local frame to the point $\bi r$}. This local frame can be taken as the Frenet-Serret frame. The surprise is that although we have postulated the existence of a free Lagrangian, the dynamical equations and the Noether analysis of the symmetries are finally independent of this free Lagrangian because the Noether constants of the motion depend only on the partial derivatives of this free Lagrangian. It is also crucial that the point $\bi r$} is moving at the speed of light. All essential observables like energy, linear and angular momentum are expressed in terms of these partial derivatives. This means that any arbitrary Poincaré invariant function $L_{0}({\bi u},{\bi a},\boldsymbol{\omega})$ of these variables will lead to the same dynamical equations and to the same definition of the fundamental observables.
Now the electron has two characterictic points $\bi q$} and $\bi r$} and therefore two different angular momenta can be defined with respect to both points. By the Noether analysis it is found that the angular momentum with respect to the CC takes the form:
${\bi S}=\left(\frac{H-{\bi u}\cdot{\bi p}}{a^{2}}\right)({\bi a}\times{\bi u}),$ (5)
perpendicular to the velocity and acceleration of the CC, and using (2) to substitute the expression of $\bi a$}, can be rewritten as
${\bi S}=-\gamma(v)m({\bi r}-{\bi q})\times{\bi u}.$ (6)
The angular momentum with respect to the CM is defined as
${\bi S}_{CM}={\bi S}+({\bi r}-{\bi q})\times{\bi p}=-{\gamma(v)}m({\bi r}-{\bi q})\times({\bi u}-{\bi v}).$ (7)
These observables represent the spin of the electron $\bi S$} and ${\bi S}_{CM}$ with respect to the CC and CM, respectively. In the center of mass frame (${\bi q}={\bi v}=0$) and they take the same value. They satisfy two different dynamical equations. In the free case $d{\bi S}/dt={\bi p}\times{\bi u}$ and $d{\bi S}_{CM}/dt=0$. The CM spin is conserved but the CC spin $\bi S$} satisfies the same dynamical equation than Dirac’s spin operator in the quantum case. It corresponds to the classical spin observable equivalent to the quantum mechanical Dirac’s spin. In the center of mass frame and for a free electron, the spin is constant and takes the value
${\bi S}=-m{\bi r}\times{\bi u},$
so that the point $\bi r$} describes a circle of constant radius $R_{0}$ at the speed of light in a plane orthogonal to the constant spin. In the center of mass frame the separation $R_{0}=|{\bi r}-{\bi q}|$ is constant and when quantizing the model takes the value $R_{0}=\hbar/2mc$, and the internal motion has an angular velocity $\omega_{0}=2mc^{2}/\hbar$, twice the frequency postulated by De Broglie. The value of the spin is $S(0)=\hbar/2$.
The angular velocity of the comoving frame $\boldsymbol{\omega}=\boldsymbol{\omega}_{p}+\boldsymbol{\omega}_{u}$, has two components, one $\boldsymbol{\omega}_{p}$ perpendicular to the plane subtended by $\bi u$} and $\bi a$} and that produces the curvature of the trajectory of the point and another $\boldsymbol{\omega}_{u}$ along the velocity $\bi u$} that gives rise to the torsion of the trajectory. They are expressed as:
$\boldsymbol{\omega}_{p}=\frac{1}{c^{2}}({\bi u}\times{\bi a}),$ (8)
$\boldsymbol{\omega}_{u}=\left[\frac{\boldsymbol{\omega}_{p}\cdot{\bi v}}{c^{2}-{\bi v}\cdot{\bi u}}\right]{\bi u}.$ (9)
There is another relation, linear in the energy and linear momentum, that depends also on the spin $\bi S$} and angular velocity:
$H={\bi p}\cdot{\bi u}+\frac{1}{c^{2}}{\bi S}\cdot\left({\bi a}\times{\bi u}\right)={\bi p}\cdot{\bi u}-{\bi S}\cdot{\boldsymbol{\omega}}.$ (10)
The last term is in fact ${\bi S}\cdot{\boldsymbol{\omega}}_{p}>0$, according to (5) and (8), because these two vectors are collinear but in opposite directions. The spin $\bi S$} is orthogonal to the velocity $\bi u$}, and therefore to $\boldsymbol{\omega}_{u}$, and this term can be rewritten as ${\bi S}\cdot{\boldsymbol{\omega}}>0$. The energy contains two terms: the first term, proportional to the linear momentum, represents the translational energy, while the second, proportional to the spin and angular velocity and to the motion of the CC (zitterbewegung), and independent of the CM velocity, represents the rotational energy $H_{rot}$ which is positive definite, because the spin $\bi S$} has the opposite direction to the angular velocity $\boldsymbol{\omega}_{p}$, and never vanishes. If we introduce in (10) the expressions of $\bi p$} in (1) and $\bi S$} in (6),
$H=\gamma(v)m{\bi v}\cdot{\bi u}+\gamma(v)m\left[({\bi r}-{\bi q})\times{\bi u}\right]\cdot\left[\frac{1}{c^{2}}{\bi u}\times{\bi a}\right],$
we get $H=\gamma(v)mc^{2}$. It is this linear expression (10) in terms of $H$ and $\bi P$}, which represents the classical equivalent of Dirac’s Hamiltonian as shown in [5].
The Classical Dirac particle has two invariant properties related to the two Casimir operators of the Poincaré grup, namely the invariant related to the four-momentum $p^{\mu}\equiv(H/c,{\bi p})$, $p^{\mu}p_{\mu}=m^{2}c^{2}$, and the invariant related to the Pauli-Lubanski four-vector $w^{\mu}=(1/2)\epsilon^{\mu\nu\sigma\lambda}p_{\nu}J_{\sigma\lambda}$, that can be written in terms of $H$, $\bi p$} and ${\bi S}_{CM}$ as:
$w^{\mu}\equiv({\bi p}\cdot{\bi S}_{CM},H{\bi S}_{CM}/c),\quad w^{\mu}w_{\mu}=-m^{2}c^{2}S(0)^{2},$ (11)
so that the two parameters, the mass $m$ and the absolute value of the spin in the center of mass frame $S(0)$, are the intrinsic properties of the electron.
Because we locate at the point $\bi r$} the total charge of the electron $e$, its motion with respect to the CM produces an instantaneous magnetic moment and an instantaneous electric dipole moment that are defined as [9]:
$\boldsymbol{\mu}_{CM}=\frac{e}{m\gamma(v)}{\bi S}_{CM}=-e({\bi r}-{\bi q})\times({\bi u}-{\bi v}),$ (12)
${\bi d}_{CM}=e({\bi r}-{\bi q}).$ (13)
In this way the magnetic moment is $\mu_{CM}=e\hbar/2m$, Bohr’s magneton and the electric dipole $d_{CM}=e\hbar/2mc$.
All physical observables $H$, $\bi p$}, the two spins $\bi S$}, ${\bi S}_{CM}$, the two components of the angular velocity $\boldsymbol{\omega}_{u}$ and $\boldsymbol{\omega}_{p}$ and the synchronous location of the CM $\bi q$} can be expressed in terms of the point $\bi r$} and their time derivatives. Also the electromagnetic properties magnetic moment $\boldsymbol{\mu}_{CM}$ and electric dipole moment ${\bi d}_{CM}$, if the particle has electric charge $e$.
In this analysis no hypothesis about the size or shape of the electron has been stated and no requirements concerning a possible charge or mass distribution. The physical properties of the electron are determined by the evolution of a single point, the center of charge. It is a system of only three degrees of freedom $\bi r$}, but they satisfy a system of ordinary differential equations of fourth order and not of second order, as it corresponds to the description of the spinless point electron. In three dimensional Euclidean space the most general continuous and differentiable curve satisfies necessarily a system of ordinary fourth order diferential equations.
The historical suggestion in some physics books that the fundamental equations of physics are second order differential equations represents a simplifying restriction. This requires that a Lagrangian description of a moving point will be a function of the point and its velocity, Euler-Lagrange equations are of second order and thus the trajectories of the point have no torsion, which is equivalent to the suppression of spin. It is the mechanical description of spinless objects and in the real world experimental physics has shown that there are no spinless elementary particles.
In the publications and preprints [8], [10], [11] we include links to some Mathematica notebooks and to Wolfram’s Community Groups where we compute numerically the interaction of the Dirac particle with uniform and oscillating electric and magnetic fields, the interaction with a polarized electromagnetic plane wave and also a Poincaré invariant interaction of two classical Dirac particles [12]. In this last analysis we find that two electrons with their spins parallel can form a metastable bound state which is not destroyed by electric fields and magnetic fields along the spin direction. Only a transversal magnetic field separates the two particles. This is not a Cooper pair because it is a spin 1 bound system of charge $2e$, separated by a distance below Compton’s wave length, smaller than the correlation distance of the Cooper pair.
Another conclusion of the formalism is that the so-called tunnel effect is not a pure quantum mechanical effect, because the usual analysis shows that a spinless point particle can never cross a potential barrier. But real particles are spinning particles so that the probability of crossing in quantum mechanics has to be compared with the analysis by using classical models of spinning particles. In the simple example [13] it is shown that even a non-relativistic spinning particle, with kinetic energy below the top of the potential, can tunnel provided its spin is orthogonal to the direction of the velocity to the barrier, while it never crosses the barrier if it is oriented in the same direction as the velocity. In any other orientation also crosses but a minimal kinetic energy below the potential barrier is required. This is known in the literature as the spin polarized tunneling effect which was used in the old magnetic hard drivers.